OpenStudy (anonymous):

i am shopping online and i find a tablet case for $16.95 and a MP3 player for $20.69 what is the price of each item after a 15% discount? Figure out the total price for both items including a 8.25% sales tax? please help me i will give medals

OpenStudy (anonymous):

total b4 taxes= 16.95+20.69 =$37.64

OpenStudy (anonymous):

with the 15% discount u are only paying 100-15=85% of the $37.64

OpenStudy (anonymous):

0.85 x 37.64=$31.994 witout taxes

OpenStudy (anonymous):

with taxes its $31.994 x 0.0825+$31.994=$34.633505 which is about $34.63

OpenStudy (anonymous):

can you help me @LadyM

OpenStudy (anonymous):

uuuuh hold on lol

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

oh boy my mind is at a dead end

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol let me ask a genius

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

i messed up because u were looking for the price of each item when i did it as a total price -.-"

OpenStudy (anonymous):

the process is similar to what i did above

OpenStudy (anonymous):